At December 31, DePaul Corporation had a $26 million balance in its deferred tax asset account and a $102 million balance in its deferred tax liability account. The balances were due to the following cumulative temporary differences :

1. Estimated warranty expense, $35 million: expense recorded in the year of sale; tax-deductible when paid (one-year warranty)

2. Depreciation expense, $180 million: straight line in the income statement; MACRS on the tax return

3. Income from installment sales of properties, $75 million: income recorded in the year of the sale; taxable when received equally over the next five years

4. Rent revenue collected in advance, $30 million; taxable in the year collected; recorded as income when the performance obligation is satisfied in the following year

Assuming DePaul will show a single noncurrent net amount in its December 31 balance sheet, indicate that amount and whether it is a net deferred tax asset or liability. The tax rate is 40%. Determine the deferred tax amounts to be reported in the December 31 balance sheet. The tax rate is 40%.   
Answer #1

Particulars

Classification

Amount

Tax rate

Deferred tax liability
(asset)

Current/Non-current

Current

Non-current

Unpaid estimated warranty expense

Current

(35)

40%

(14)

Depreciation

Non-current

180

40%

72

Income from instalment sales

Current

15

40%

6

Income from instalment sales

Non-current

60

40%

24

Unearned rent revenue

Current

(30)

40%

(12)

Net current liability (asset)

(20)

Net non-current liability (asset)

96

This is shown in balance sheet as

Current Assets

       Deferred taxes

(20)

Long term Liabilities

       Deferred taxes

96

Before offsetting assets and liabilities within the current and non-current categories, the total deferred tax assets is 14+12 = $ 26 million and deferred tax liability is 6+72+24 = $ 102 million

Notes :

Current portion of income from instalment sales = 75/5 = $ 15 million

The balance amount is non- current portion as = 75-15 = $ 60 million

Kindly upvote